calculate earth's mass given the acceleration due to gravity at the North pole is 9.830 m/s² and the radius of the earth is 6371km from pole to pole.

1
Expert's answer
2022-02-28T10:41:16-0500
"g=G\frac{M}{R^2}"

"M=gR^2/G\\
=9.8*(6371*10^3)^2/(6.67*10^{-11})\\
=5.96*10^{24}\:\rm kg"